# R-Zero: Self-Evolving Reasoning LLM from Zero Data

### v3

@@ 5 Analysis
In this section, we conduct a series of in-depth analyses to better understand the behavior and effectiveness of our R-Zero framework. To ensure consistency, all analytical experiments presented here were conducted on the Qwen3-4B-Base model, unless explicitly stated otherwise.
@@ 5.1 Ablation Study
Table 3: Ablation study results on the Qwen3-4B-Base model. w/o RL-Challenger: Disables GRPO training for the Challenger. w/o Filtering: Disables the difficulty-based curriculum filtering. w/o Rep. Penalty: Removes the repetition penalty from the Challenger’s reward.
@@
   $\vdash$ w/o RL-Challenger 44.36 26.32
   $\vdash$ w/o Rep. Penalty 45.76 27.56
   $\vdash$ w/o Filtering 47.35 24.26
To isolate the contribution of each key component within our R-Zero framework, we conduct a comprehensive ablation study on the Qwen3-4B-Base model. We specifically investigate the importance of three critical modules by disabling them one at a time and observing the impact on performance. The results are summarized in Table 3.
As shown in the table, removing any core components leads to a significant degradation in performance. The largest drop occurs when we disable the Challenger’s reinforcement learning (w/o RL-Challenger), with the Math and General average scores decreasing by 3.7 and 4.1 points, respectively. This result highlighting the importance of our co-evolutionary curriculum generation process. Similarly, removing the Repetition Penalty also harms performance, indicating that generating a diverse set of questions is crucial for effective Solver training.
Finally, disabling the Task Filtering module results in a notable performance drop, particularly on the general-domain average, which falls by over 6 points. As discussed in Section 3.3, this filtering serves a dual purpose: it calibrates the curriculum’s difficulty and acts as an implicit quality control mechanism by removing questions with low answer consistency. Without this filter, the Solver is trained on a noisy and poorly curated dataset that likely includes ambiguous or ill-posed questions, which harms its ability to learn robustly.
@@ 5.2 Evolution of Question Difficulty and Data Accuracy
Base Model Solver (Iter 1) Solver (Iter 2) Solver (Iter 3) Pseudo-Label Acc.
$\mathcal{D}_{\text{Iter 1}}$ 48.0 59.0 57.0 61.0 79.0%
$\mathcal{D}_{\text{Iter 2}}$ 52.5 53.0 51.5 53.5 69.0%
$\mathcal{D}_{\text{Iter 3}}$ 44.0 47.0 45.0 50.5 63.0%
To understand the co-evolutionary dynamic, we analyzed how the Challenger’s generated questions and their corresponding pseudo-labels change across iterations. We sampled 200 questions from the Challenger’s policy after each of the first three training iterations, creating three distinct test sets: $\mathcal{D}_{\text{Iter 1}}$, $\mathcal{D}_{\text{Iter 2}}$, and $\mathcal{D}_{\text{Iter 3}}$. For this analysis, we assumed the external oracle model, GPT-4o, to be a perfect annotator, providing the ground truth answers for all generated questions.
@@
This analysis reveals a multi-faceted dynamic. The first finding is that the questions generated by the Challenger become progressively more difficult. This is directly evidenced by evaluating a fixed model against the evolving question sets. For instance, the performance of the static Solver (Iter 1), when measured against the consistent GPT-4o ground truth, drops from 59.0% on Iteration 1 questions to 47.0% on Iteration 3 questions. This confirms that the Challenger is successfully increasing the intrinsic difficulty of its curriculum. The second finding, revealed by the highlighted column, pertains to the true accuracy of the self-generated dataset. Unfortunately, while the accuracy of the pseudo-labels is initially high at 79.0%, it systematically drops to 63.0% by the third iteration. This trend indicates that as the system generates more difficult problems, the Solver’s majority vote becomes a less reliable source for ground truth. This decline in data quality is a critical trade-off and a potential bottleneck for the framework’s ultimate performance.
Finally, despite this drop in absolute label accuracy, the framework’s internal reward mechanism functions precisely as designed. The scores on the table’s diagonal show how each Solver performs on questions from its contemporary Challenger. The Solver (Iter 2) achieves 51.5% and the Solver (Iter 3) achieves 50.5% on their respective question sets. This demonstrates that the Challenger successfully calibrates the question difficulty to match the Solver’s evolving capabilities, consistently targeting the 50% success rate that our reward function incentivizes.
5.3 Synergy with Supervised Data
We first establish a supervised baseline by fine-tuning the base model directly on the labeled data. For this process, we employ GRPO, an approach similar to Zero-RL (Zeng et al., 2025).
We then apply our R-Zero framework, where at the end of each co-evolutionary iteration, the resulting checkpoint is also fine-tuned on the same labeled dataset. The results show that our method provides significant additional gains. As highlighted in Figure 3, this represents a gain of +2.35 points over the direct training baseline.
5.4 Iteration Scaling
Previous results demonstrate that  R-Zero generally enhance the Solver’s capabilities across iterations. A closer inspection, however, reveals that the improvement is not consistent, with performance on certain challenging benchmarks degrading in later iterations. This raises a critical question about the long-term stability of our self-improvement loop: what are the limits of this process, and what causes this eventual performance degradation? In this section, we conduct a dedicated analysis to investigate these iteration scaling dynamics, aiming to diagnose the underlying cause of this instability.
5.4.1 The Inevitability of Collapse: An Empirical Analysis
For instance, the smallest 0.6B model reaches its peak performance as early as the first iteration (Iter 1), after which its capabilities begin to decline. In contrast, the largest 4B model sustains its upward trajectory for three full iterations, only experiencing a sharp drop at Iter 4. This pattern strongly suggests that while larger model capacity can delay the negative effects, it does not prevent them. This eventual collapse points to an inherent instability or limitation within our current self-improvement framework, highlighting a critical area for future investigation.
5.4.2 Beyond Label Noise: Unpacking the Roots of Instability
Iteration Model Size
0.6B 1.7B 4B
Iter 1 70.6 69.4 71.0
Iter 2 53.4 55.2 56.2
Iter 3 50.8 52.2 48.8
Iter 4 44.0 45.2 42.2
Although the degradation of pseudo-label accuracy is a consistent trend across iterations, our analysis suggests this is not the primary, nor even the sole, driver of the eventual performance collapse. Table 5 presents the pseudo-label data quality for each model at the onset of its performance collapse. Intriguingly, there appears to be no universal accuracy threshold that triggers this degradation. For instance, the 0.6B model begins to decline when data accuracy is still as high as 70.6% (Iter 1), whereas the 4B model tolerates an accuracy as low as 48.8% (Iter 3) before its performance drops.
This suggests that the absolute percentage of label noise is not the sole determinant of instability. Another potential, and perhaps more fundamental, reason is a form of model collapse that can be introduced when training exclusively on self-synthesized data (Tan et al., 2024a; Shumailov et al., 2024; Dohmatob et al., 2024b; Zhou et al., 2025b; Seddik et al., 2024; Dohmatob et al., 2024a; Briesch et al., 2023; Zheng et al., 2025a). A model can enter a degenerative feedback loop, suffering from a loss of diversity or an amplification of its own biases, which presents a significant challenge.
5.5 Parameter Sharing Between Challenger and Solver
R-Zero (ours) Single-R-Zero
Iter 1 48.06 71.0 47.31 63.4
Iter 2 48.44 56.2 46.95 46.6
Iter 3 49.07 48.8 45.57 32.6
Iter 4 46.52 42.2 43.89 33.8
@@ 6.1 Label-Free Reinforcement Learning
A significant trend in recent research is Label-Free Reinforcement Learning, which aims to improve LLM reasoning without human-annotated data. Many such methods use the model’s own outputs as a reward signal. This includes leveraging sequence-level confidence (Li et al., 2025a; Prabhudesai et al., 2025), the consistency of answers derived from varied reasoning paths (Zhang et al., 2025a; Zuo et al., 2025; Zhang et al., 2025b), minimizing the output entropy (Agarwal et al., 2025; Cheng et al., 2025), or even random (Shao et al., 2025) or negative reward (Zhu et al., 2025). These signals are often used within self-training loops where models fine-tune on their own most plausible solutions (Shafayat et al., 2025; Zhao et al., 2025b). While these methods all rely on a pre-existing set of unlabeled problems, R-Zero removes the need for any seed dataset.
@@ 6.2 Self-Play in Large Language Models
The paradigm of self-play, where models take on dual roles to create a self-improvement loop (Chen et al., 2024), has recently been adapted to improve language models without human data. This approach has been particularly fruitful in verifiable domains like code generation, where a “Coder” agent’s program is verified by a “Tester” agent’s unit tests (Lin et al., 2025; Wang et al., 2025a; Pourcel et al., 2025). More advanced frameworks push autonomy further by learning to generate the problems themselves, creating an adaptive curriculum from a small seed of examples or from scratch (Zhao et al., 2025a; Li et al., 2025c; Zhou et al., 2025a; Fang et al., 2025). Our work distinguishes itself by extending this paradigm to general reasoning domains that lack such verifiable environments, instead learning from a reward signal derived from the model’s own internal consistency.
@@ 6.3 Reinforcement Learning with Verifiable Rewards (RLVR)
Reinforcement Learning with Verifiable Rewards (RLVR) has been widely adopted as a versatile paradigm for enhancing LLMs across a multitude of tasks. Its effectiveness is demonstrated in diverse applications such as relation extraction (Dai et al., 2025), interactive GUI navigation (Shi et al., 2025b) and search-engine utilization (Jin et al., 2025). While early implementations relied on rule-based verifiers, recent work has begun to explore more sophisticated, model-based verifiers (Ma et al., 2025; Li et al., 2025b; 2024).
@@ 7 Conclusion and Future Work
In this paper, we introduced R-Zero, a fully autonomous self-evolving framework that overcomes data dependency by having a Challenger and Solver co-evolve to create a self-generating curriculum. Our experiments demonstrate that R-Zero significantly improves LLM’s reasoning capability on multiple domains. Future work could further focus on improving efficiency, exploring more robust labeling techniques, and expanding R-Zero to new domains. It is crucial to note, however, that the core mechanism of R-Zero is currently suited for domains where correctness can be objectively determined. Extending this self-evolutionary paradigm to open-ended generative tasks, such as creative writing or dialogue, where evaluation is subjective, remains a significant hurdle for future research. We believe R-Zero is a significant step towards creating truly self-evolving LLMs.
Appendix A Experiment Details
A.1 Training Hyperparameter
@@ A.3 GPT-4o Judge Prompt
A.4 Repetition Penalty Implementation

### v4

@@ 4 Analysis
In this section, we conduct a series of in-depth analyses to better understand the behavior and effectiveness of our R-Zero framework. To ensure consistency, all analytical experiments presented here are conducted on the Qwen3-4B-Base model, unless explicitly stated otherwise. Some additional analyses are shown in Appendix D.
@@ 4.1 Ablation Study
Table 3: Ablation results on Qwen3-4B-Base. w/o Filtering: Disables the difficulty-based curriculum filtering. w/o Rep. Penalty: Removes the repetition penalty from the Challenger’s reward.
@@
   $\vdash$ w/o Rep. Penalty 45.76 28.73
   $\vdash$ w/o Filtering 47.35 26.69
To isolate the contribution of each key component within our R-Zero framework, we conduct a comprehensive ablation study on the Qwen3-4B-Base model. We specifically investigate the importance of two critical modules by disabling them one at a time and observing the impact on performance. The results are summarized in Table 3.
As shown in the table, removing any core components leads to a significant degradation in performance. Removing the Repetition Penalty harms performance, indicating that generating a diverse set of questions is crucial for effective Solver training.
Finally, disabling the Task Filtering module results in a notable performance drop, particularly on the general-domain average, which falls by over 6 points. As discussed in Section 2.3, this filtering serves a dual purpose: it calibrates the curriculum’s difficulty and acts as an implicit quality control mechanism by removing questions with low answer consistency. Without this filter, the Solver is trained on noisy data that likely includes ambiguous or ill-posed questions, which harms its ability to learn robustly.
4.2 Iteration Scaling
Previous results demonstrate that  R-Zero enhances the Solver’s capabilities. This raises a critical question about the long-term stability of our self-improvement loop: what are the limits of this process, and whether the eventual performance degradate? In this section, we conduct an analysis to investigate these iteration scaling dynamics, aiming to diagnose the underlying cause of this instability.
For instance, the smallest 0.6B model reaches its peak performance as early as the first iteration (Step 15), after which its capabilities begin to decline. In contrast, the largest 4B model sustains its upward trajectory for three full iterations, only experiencing a sharp drop at Step 60. This pattern strongly suggests that while larger model capacity can delay the negative effects, it does not prevent them. This eventual collapse points to an inherent instability or limitation within our current self-improvement framework, highlighting a critical area for future investigation. We present some additional analysis results for this in Appendix E.
4.3 Synergy with Supervised Data
We first establish a supervised baseline by fine-tuning the base model directly on the labeled data. For this process, we also employ GRPO.
We then apply our R-Zero framework, where at the end of each co-evolutionary iteration, the resulting checkpoint is also fine-tuned on the same labeled dataset. The results show that our method provides significant additional gains. As highlighted in Figure 4, this represents a gain of +2.35 points over the human-label-only baseline for 4B model.
We also studied the scenario where human-labeled datasets are mixed into the R-Zero generated dataset for training. We found that this concurrent training approach outperforms using either R-Zero dataset or humanlabel dataset alone, but it does not surpass the sequential strategy of first applying R-Zero and then performing SFT. We hypothesize that mixing the labeled data during R-Zero training may dilute the high-quality signal while partially mitigating noise. Performing R-Zero first allows the model to acquire reasoning ability before leveraging the high-quality human-labeled data, which appears to be the most effective strategy. The precise underlying reasons require further investigation.
Dataset AMC Minerva MATH GSM8K Olympiad AIME25 AIME24 SuperGPQA MMLU-Pro BBEH
Human 57.97 55.15 80.8 92.04 48 9.58 10.31 29.49 57.03 9.71
R-Zero 57.27 52.94 79.6 92.12 44.59 9.6 13.4 27.55 55.47 10.42
R-Zero+Human 57.9 53.2 81.2 92.2 47.7 10.32 14.67 29.4 58.2 11.8
Table 4: Performance comparison across benchmarks when training with human-labeled data, R-zero self-generated data, and a mixed training strategy.
@@ 4.4 Evolution of Question Difficulty and Data Accuracy
Base Model Solver (step 15) Solver (step 30) Solver (step 45) Pseudo-Label Acc.
$\mathcal{D}_{\text{Step 15}}$ 48.0 59.0 57.0 61.0 79.0%
$\mathcal{D}_{\text{Step 30}}$ 52.5 53.0 51.5 53.5 69.0%
$\mathcal{D}_{\text{Step 45}}$ 44.0 47.0 45.0 50.5 63.0%
To understand the co-evolutionary dynamic, we analyzed how the Challenger’s generated questions and their corresponding pseudo-labels change across iterations. We sampled 200 questions from the Challenger’s policy after each of the first three training iterations, creating three distinct test sets: $\mathcal{D}_{\text{step 15}}$, $\mathcal{D}_{\text{step 30}}$, and $\mathcal{D}_{\text{step 45}}$. For this analysis, we assumed the external oracle model, GPT-4o, to be a perfect annotator, providing the ground truth answers for all generated questions.
@@
This analysis reveals a multi-faceted dynamic. The first finding is that the questions generated by the Challenger become progressively more difficult. This is directly evidenced by evaluating a fixed model against the evolving question sets. For instance, the performance of the static Solver (Step 15), when measured against the consistent GPT-4o ground truth, drops from 59.0% on the questions for the Step 15 training to 47.0% on the questions for the Step 45. This confirms that the Challenger is successfully increasing the intrinsic difficulty of its curriculum.
The second finding, revealed by the highlighted column, pertains to the true accuracy of the self-generated dataset. Unfortunately, while the accuracy of the pseudo-labels is initially high at 79.0%, it systematically drops to 63.0% by the third iteration. This trend indicates that as the system generates more difficult problems, the Solver’s majority vote becomes a less reliable source for ground truth. This decline in data quality is a critical trade-off and a potential bottleneck for the framework’s ultimate performance.
@@ 5.1 Label-Free Reinforcement Learning
A significant trend in recent research is Label-Free Reinforcement Learning, which aims to improve LLM reasoning without human-annotated data. Many such methods use the model’s own outputs as a reward signal. This includes leveraging sequence-level confidence (Li2025ConfidenceIA; Prabhudesai2025MaximizingCA), the consistency of answers derived from varied reasoning paths (Zhang2025ConsistentPL; Zuo2025TTRLTR; Zhang2025RightQI; zhou2025evolving; prasad2024self), minimizing the output entropy (Agarwal2025TheUE; Cheng2025ReasoningWE), or even random (shao2025spuriousrewardsrethinkingtraining) or negative reward (Zhu2025TheSE). These signals are often used within self-training loops where models fine-tune on their own most plausible solutions (Shafayat2025CanLR; Zhao2025LearningTR). While these methods all rely on a pre-existing set of unlabeled problems, R-Zero removes the need for any seed dataset.
@@ 5.2 Self-Play in Large Language Models
The paradigm of self-play, where models take on dual roles to create a self-improvement loop (chen2024self; zhang2024survey), has recently been adapted to improve language models without human data. This approach has been particularly fruitful in verifiable domains like code generation, where a “Coder” agent’s program is verified by a “Tester” agent’s unit tests (Lin2025LearningTS; Wang2025CoEvolvingLC; Pourcel2025SelfImprovingLM; jiang2025verse). More advanced frameworks push autonomy further by learning to generate the problems themselves, creating an adaptive curriculum from a small seed of examples or from scratch (Zhao2025AbsoluteZR; li2025videohalluevaluatingmitigatingmultimodal; Zhou2025SelfChallengingLM; Fang2025SeRLSR). Our work distinguishes itself by extending this paradigm to general reasoning domains that lack such verifiable environments, like coding tasks.
@@ 5.3 Reinforcement Learning with Verifiable Rewards (RLVR)
Reinforcement Learning with Verifiable Rewards has been widely adopted as a versatile paradigm for enhancing LLMs across a multitude of tasks (li2025self; DeepSeekAI2025DeepSeekR1IR; Shao2024DeepSeekMathPT). Its effectiveness is demonstrated in diverse applications such as relation extraction (dai2025r1), interactive GUI navigation (Shi2025MobileGUIRLAM) and search-engine utilization (Jin2025SearchR1TL). While early implementations relied on rule-based verifiers, recent work has begun to explore more sophisticated, model-based verifiers (Ma2025GeneralReasonerAL; li2025semanticallyawarerewardsopenendedr1; li2024fg).
6 Limitation
While R-Zero demonstrates strong improvements in reasoning performance across multiple domains, several limitations remain. First, the core mechanism of R-Zero relies on domains where correctness can be objectively verified. The Challenger–Solver co-evolution process depends on clear, deterministic evaluation signals to produce reliable training feedback. Consequently, applying R-Zero to open-ended or subjective tasks, such as creative writing, dialogue, or preference-driven generation, remains difficult, as these tasks lack unambiguous correctness criteria. In addition, R-Zero currently employs specific labeling and verification strategies that may not generalize to all task types. Developing more robust and broadly applicable labeling mechanisms would further expand the range of domains where R-Zero can be effectively applied.
@@ 7 Conclusion and Future Work
In this paper, we introduced R-Zero, a fully autonomous self-evolving framework that overcomes data dependency by having a Challenger and Solver co-evolve to create a self-generating curriculum. Our experiments demonstrate that R-Zero significantly improves LLM’s reasoning capability in multiple domains. Future work could further focus on improving efficiency, exploring more robust labeling techniques, and expanding R-Zero to new domains. Extending this self-evolutionary paradigm to open-ended generative tasks, such as creative writing or dialogue, where evaluation is subjective, remains a significant hurdle for future research.
Reproducibility Statement
To ensure the reproducibility of our research, we provide detailed information regarding our data and experimental setup. All datasets used in this work are publicly available; we provide details on data sources, any postprocess steps in Appendix B. A comprehensive list of all hyperparameters for each experiment can be found in a table in Appendix B and Sec. 3.
Appendix
Appendix A The Use of Large Language Models (LLMs)
We acknowledge the use of large language models (LLMs) as assistive tools in this research, with usage limited to refining grammar and improving language clarity in the manuscript, writing utility scripts for data preprocessing and postprocessing, and debugging; all outputs from these models were meticulously reviewed, revised, and verified by the authors, who retain full responsibility for all content presented in this paper.
Appendix B Experiment Details
B.1 Training Hyperparameter
@@ B.3 GPT-4o Judge Prompt
B.4 Repetition Penalty Implementation
@@ 3. Final Penalty Calculation
Appendix C Evaluation Benchmark
Mathematical Reasoning.
We use seven challenging benchmarks: AMC, Minerva (Lewkowycz2022SolvingQR), MATH-500 (Hendrycks2021MeasuringMP), GSM8K (Cobbe2021TrainingVT), Olympiad-Bench (He2024OlympiadBenchAC), AIME-2024, and AIME-2025. For these tasks, where answers can be complex, we employ GPT-4o as a programmatic judge to semantically verify the correctness of the final answer against the ground truth (Zhao2025OneTT). For the difficult AMC and AIME benchmarks, we report the mean@32 metric. For all other math benchmarks, we report accuracy based on greedy decoding.
General Domain Reasoning.
To test for the generalization of reasoning ability, we evaluate the following challenging benchmarks:
  • MMLU-Pro (Wang2024MMLUProAM): An enhanced version of the MMLU (Hendrycks2020MeasuringMM) benchmark, featuring a more challenging suite of multi-task questions designed to provide a stricter evaluation of language model capabilities.
  • SuperGPQA (Team2025SuperGPQASL): A large-scale benchmark focused on graduate-level reasoning. It comprises questions across 285 distinct disciplines that have been verified as unsearchable on the web, thereby isolating true reasoning ability from simple knowledge recall.
  • BBEH (kazemi2025BIGBenchEH): This benchmark builds upon the foundation of BIG-Bench Hard (Suzgun2022ChallengingBT) by incorporating a new selection of tasks specifically engineered to be more difficult, thus providing a more accurate measure of complex reasoning skills.
For this category, we follow the experimental setup, prompts, and evaluation codes from (Ma2025GeneralReasonerAL), reporting Exact Match (EM) accuracy obtained via greedy decoding.
Appendix D Parameter Sharing Between Challenger and Solver
R-Zero (ours) Single-R-Zero
Step 15 48.06 71.0 47.31 63.4
Step 30 48.44 56.2 46.95 46.6
Step 45 49.12 48.8 45.57 32.6
Step 60 46.52 42.2 43.89 33.8
Appendix E Beyond Label Noise: Unpacking the Roots of Instability
Step Model Size
0.6B 1.7B 4B
Step 15 70.6 69.4 71.0
Step 30 53.4 55.2 56.2
Step 45 50.8 52.2 48.8
Step 60 44.0 45.2 42.2
Although the degradation of pseudo-label accuracy is a consistent trend across iterations, our analysis suggests this is not the primary, nor even the sole, driver of the eventual performance collapse. Table 7 presents the pseudo-label data quality for each model at the onset of its performance collapse. Intriguingly, there appears to be no universal accuracy threshold that triggers this degradation. For instance, the 0.6B model begins to decline when data accuracy is still as high as 70.6% (Step 15), whereas the 4B model tolerates an accuracy as low as 48.8% (Step 45) before its performance drops.
This suggests that the absolute percentage of label noise is not the sole determinant of instability. Another potential, and perhaps more fundamental, reason is a form of model collapse that can be introduced when training exclusively on self-synthesized data (tan2024large; Shumailov2024AIMC; Dohmatob2024ATO; zhou2025dissecting; Seddik2024HowBI; Dohmatob2024StrongMC; Briesch2023LargeLM; zheng2025learning). A model can enter a degenerative feedback loop, suffering from a loss of diversity or an amplification of its own biases, which presents a significant challenge.
Appendix F Theoretical Analysis
In this section, we provide a theoretical motivation for our uncertainty reward function, $r_{\text{uncertainty}}\propto 1-2|\hat{p}(x;S_{\phi})-\tfrac{1}{2}|$, which is maximized when the Solver’s success probability, $\hat{p}$, is 50%. Our analysis is grounded in recent work that formally establishes that the most efficient training occurs when a learner is exposed to tasks at the frontier of its capabilities (Shi2025EfficientRF; Bae2025OnlineDF).
$\mathbb{D}_{KL}(S_{\phi}||S^{*})\geq\frac{\hat{p}(1-\hat{p})}{2\beta^{2}}$
where $\beta$ is the temperature parameter that controls entropy regularization. The right-hand side of the inequality, which is proportional to the reward variance, is maximized precisely when $\hat{p}=0.5$. Therefore, by designing the Challenger’s reward to incentivize questions that push the current Solver towards this point of maximum uncertainty, our framework is theoretically motivated to generate a maximally efficient curriculum in each iteration of the co-evolutionary process.
Appendix G Algorithm of R-Zero
Here, we provide the pseudocode for our method.
Input:Initial models $Q_{\theta},S_{\phi}$; Group size $G$; Solver samples $m$; Dataset size NN; Filtering threshold $\delta$; Repetition threshold $\tau_{\text{BLEU}}$. Output:Evolved models $Q_{\theta}$ and $S_{\phi}$.
for each self-play iteration do
  // --- Phase 1: Challenger Training (Sec 2.2) ---
  for each Challenger training step do
     Sample question group $\{x_{i}\}_{i=1}^{G}\sim Q_{\theta}(\cdot)$;
     for each question $x_{i}$ do
        if FormatCheck($x_{i}$) is invalid (e.g., missing tags) then
           $r_{i}\leftarrow 0$;
        end if
       else
           Sample $m$ answers $\{y_{j}\}_{j=1}^{m}\sim S_{\phi}(\cdot\mid x_{i})$;
           Get pseudo-label $\tilde{y}_{i}\leftarrow\text{MajorityVote}(\{y_{j}\}_{j=1}^{m})$;
           Compute correctness $\hat{p}_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}\mathbbm{1}\{y_{j}=\tilde{y}_{i}\}$;
           Compute uncertainty reward $r_{\text{uncertainty}}\leftarrow 1-2|\hat{p}_{i}-\tfrac{1}{2}|$;
           Compute repetition penalty $r_{\text{rep}}(x_{i})$ via BLEU clustering (Sec 2.2);
           Final reward: $r_{i}\leftarrow\max\bigl(0,r_{\text{uncertainty}}-r_{\text{rep}}(x_{i})\bigr)$;
        end if
     end for
    Update $Q_{\theta}$ via GRPO using rewards $\{r_{i}\}_{i=1}^{G}$;
  end for
 // --- Phase 2: Solver Dataset Construction (Sec 2.3) ---
  Initialize curated dataset $\mathcal{S}\leftarrow\emptyset$;
  Sample NN candidate questions $\{x_{k}\}_{k=1}^{N}\sim Q_{\theta}(\cdot)$;
  for each candidate $x_{k}$ do
     Sample $m$ answers $\{y_{j}\}_{j=1}^{m}\sim S_{\phi}(\cdot\mid x_{k})$;
     Get pseudo-label $\tilde{y}_{k}\leftarrow\text{MajorityVote}(\{y_{j}\})$;
     Compute correctness $\hat{p}_{k}\leftarrow\frac{1}{m}\sum_{j=1}^{m}\mathbbm{1}\{y_{j}=\tilde{y}_{k}\}$;
     if $|\hat{p}_{k}-\tfrac{1}{2}|\leq\delta$ then
        Add $(x_{k},\tilde{y}_{k})$ to $\mathcal{S}$;
     end if
  end for
 // --- Phase 3: Solver Training (Sec 2.4) ---
  for each minibatch $(x,\tilde{y})\in\mathcal{S}$ do
     Sample $G$ answers $\{y_{j}\}_{j=1}^{G}\sim S_{\phi}(\cdot\mid x)$;
[…]
     Update $S_{\phi}$ via GRPO using rewards $\{r_{j}\}_{j=1}^{G}$;
  end for
end for
Algorithm 1 R-Zero: Self-Evolving Challenger-Solver Framework
Appendix H Generalizing Beyond the Math Domain
To further evaluate the generality of our proposed approach, we conduct an additional experiment in which the model is no longer instructed to generate exclusively mathematical problems during data construction. Instead, we remove the math-specific constraint from the prompting stage, allowing the model to produce a broader mixture of questions, including commonsense, logical reasoning, and other non-mathematical categories. This setup enables us to examine whether the improvements observed in earlier experiments are merely a consequence of alignment with math-heavy benchmarks or whether our method can maintain (or even enhance) performance under a more domain-general training regime.
The motivation behind introducing this experiment stems from concerns that many mathematical benchmarks contain well-structured and heavily studied problem formats. If the training data is restricted to similar mathematical tasks, it may artificially inflate performance gains while providing limited evidence regarding the method’s real-world applicability. By diversifying the data generation process, we aim to assess whether the model can benefit from richer and more heterogeneous supervision signals.
The experimental results, summarized in Table 8, demonstrate that when the base model has sufficient capacity (e.g., the 8B variant), training on more general data leads to further improvements in both math and non-math benchmarks. Notably, the “R-Zero (remove math prompt)” setting exhibits the highest performance on the general benchmarks for the 8B model, suggesting that our approach effectively transfers to broader reasoning tasks beyond pure mathematics.
Table 8: Results of different models and prompting strategies on mathematics and reasoning benchmarks.
AMC Minerva MATH GSM8K Olympiad AIME25 AIME24 SuperGPQA MMLU-Pro BBEH
Qwen3-4B 45.7 38.24 68.2 87.79 41.04 10.3 6.7 20.88 37.38 7.57
R-Zero 57.27 52.94 79.6 92.12 44.59 9.6 13.4 27.55 55.47 10.42
R-Zero (no math) 56.27 52.56 79.2 92.56 43.56 9.9 13.2 28.14 54.75 10.31
Qwen3-8B 51.95 50 78 89.08 44.74 12.1 14.6 28.33 51.8 8.63
R-Zero 61.67 60.66 82 94.09 48.89 13.3 15.4 31.38 61.53 10.60
R-Zero (no math) 65.53 61.19 81.6 93.89 49.24 13.2 15.2 32.29 61.83 10.88
These findings provide additional evidence that the proposed method is not limited to mathematical domains; rather, it scales effectively with model capacity and supports improved reasoning performance under more diverse training conditions.
Appendix I Preliminaries
I.1 Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) (Shao2024DeepSeekMathPT) is a reinforcement learning algorithm for fine-tuning policy LLMs $\pi_{\theta}$ without a separately learned value function. Its key idea is to normalize rewards within a group of responses sampled from the same prompt, thereby stabilizing optimization.
Setup.
Given a query $\mathbf{q}$, the old policy $\pi_{\theta_{\text{old}}}$ generates $G$ candidate responses $\{\mathbf{o}_{1},\dots,\mathbf{o}_{G}\}$. Each response $\mathbf{o}_{i}$ is assigned a scalar reward $R(\mathbf{q},\mathbf{o}_{i})$. Group-wise z-score normalization is then applied to obtain an advantage shared across the tokens of the response:
$\hat{A}_{i,t}=\frac{R(\mathbf{q},\mathbf{o}_{i})-\operatorname{mean}\!\bigl(\{R(\mathbf{q},\mathbf{o}_{1}),\dots,R(\mathbf{q},\mathbf{o}_{G})\}\bigr)}{\operatorname{std}\!\bigl(\{R(\mathbf{q},\mathbf{o}_{1}),\dots,R(\mathbf{q},\mathbf{o}_{G})\}\bigr)+\varepsilon_{\text{norm}}},$
where $\varepsilon_{\text{norm}}$ is a small constant for numerical stability. The same normalized advantage $\hat{A}_{i,t}$ is applied to all tokens of response $\mathbf{o}_{i}$.
Policy Update.
Let $r_{\theta}(o_{i,t})\;=\;\frac{\pi_{\theta}(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid\mathbf{q},\mathbf{o}_{i,<t})}.$ The policy is updated with a clipped surrogate objective, similar to PPO, combined with a KL-divergence penalty to constrain policy drift:
$\mathcal{L}_{\text{GRPO}}(\theta)=-\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|\mathbf{o}_{i}|}\sum_{t=1}^{|\mathbf{o}_{i}|}\min\!\Bigl(r_{\theta}(o_{i,t})\,\hat{A}_{i,t},\;\operatorname{clip}\bigl(r_{\theta}(o_{i,t}),\,1-\epsilon,\,1+\epsilon\bigr)\,\hat{A}_{i,t}\Bigr)\;+\;\beta\,\operatorname{KL}\!\bigl(\pi_{\theta}(\mathbf{q})\,\|\,\pi_{\theta_{\text{old}}}(\mathbf{q})\bigr)$
Maximizing the negative of this loss encourages the policy to increase the probability of tokens contributing to responses with positive relative advantages, while clipping prevents overly aggressive updates. The KL penalty, weighted by $\beta$, further stabilizes training by preventing the new policy from drifting too far from the old one.
I.2 Reinforcement Learning with Verifiable Rewards
Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert2024TLU3P) is a paradigm for fine-tuning models in domains where response quality can be deterministically verified. RLVR relies on a rule-based verifier $v:\mathcal{X}\to\{0,1\}$ that assigns a binary reward to each response $x_{i}$: